# Vertex-based Graph Neural Solver and its Application to Linear Elasticity Equations

### v1

@@ 3 Datasets
For the 2D case, the stress tensor $\boldsymbol{\sigma}$ and strain tensor $\boldsymbol{\varepsilon}$ (using Voigt notation) are expressed as:
$\boldsymbol{\sigma}=\begin{bmatrix}\sigma_{xx}\\[2.0pt] \sigma_{yy}\\[2.0pt] \sigma_{xy}\end{bmatrix},\qquad\boldsymbol{\varepsilon}=\begin{bmatrix}\varepsilon_{xx}\\[2.0pt] \varepsilon_{yy}\\[2.0pt] 2\varepsilon_{xy}\end{bmatrix}$. (15)
The corresponding isotropic stiffness matrix $\boldsymbol{C}$ takes the following form:
$\boldsymbol{C}=\begin{bmatrix}\lambda+2\mu&\lambda&0\\ \lambda&\lambda+2\mu&0\\ 0&0&\mu\end{bmatrix}$,
where the Lamé constants are defined as:
@@
We consider the Poisson’s ratio $\nu\in[0,0.5)$ to be constant, while the Young’s modulus $E(\boldsymbol{x})$ follows a log-normal Gaussian random field (GRF) distribution [42] to simulate spatially heterogeneous materials:
$E(\boldsymbol{x})=\alpha_{m}\cdot\exp(w(\boldsymbol{x}))+\beta_{m},$ (17)
where $\alpha_{m}=10^{8}$ and $\beta_{m}=100$. The random field $w(\boldsymbol{x})\sim\mathcal{N}(0,L_{\Delta}^{-2})$, with its covariance operator $L_{\Delta}^{-2}$ defined by the following second-order elliptic operator:
@@
The computational domain is $\Omega=[0,1]^{2}$. The boundary conditions are defined as follows: the left boundary $\Gamma_{1}=\{\boldsymbol{x}\in\partial\Omega\mid x_{1}=0\}$ is fixed (Dirichlet), and a traction force $\boldsymbol{t}=(10^{6},0)^{\top}$ is applied to the right boundary $\Gamma_{2}=\{\boldsymbol{x}\in\partial\Omega\mid x_{1}=1\}$ (Neumann). The body force is set to $\boldsymbol{f}=(0,0)^{\top}$, and the Poisson’s ratio is fixed at $\nu=0.4$. Figure 1 illustrates the mesh discretization of the domain, a realization of the Young’s modulus field, and the corresponding displacement solution. By sampling the random field $E(\boldsymbol{x})$ $M$ times and discretizing the resulting PDEs, we obtain $M$ training samples.
@@
For the 3D case, the stress and strain tensors are expressed as:
$\boldsymbol{\sigma}=[\sigma_{xx},\sigma_{yy},\sigma_{zz},\sigma_{yz},\sigma_{xz},\sigma_{xy}]^{\top},\quad\boldsymbol{\varepsilon}=[\varepsilon_{xx},\varepsilon_{yy},\varepsilon_{zz},2\varepsilon_{yz},2\varepsilon_{xz},2\varepsilon_{xy}]^{\top}.$ (19)
The corresponding stiffness matrix $\boldsymbol{C}$ is given by:
$\boldsymbol{C}=\begin{bmatrix}\lambda+2\mu&\lambda&\lambda&0&0&0\\ \lambda&\lambda+2\mu&\lambda&0&0&0\\ \lambda&\lambda&\lambda+2\mu&0&0&0\\ 0&0&0&\mu&0&0\\ 0&0&0&0&\mu&0\\ 0&0&0&0&0&\mu\end{bmatrix}$.
The generation of Young’s modulus follows the same procedure as in Data-1. The computational domain is shown in Figure 2. The boundary conditions are: the left boundary $\Gamma_{1}=\{\boldsymbol{x}\in\partial\Omega\mid x_{1}=0\}$ is fixed, and a traction force $\boldsymbol{t}=(10^{6},0,0)^{\top}$ is applied to the right boundary $\Gamma_{2}=\{\boldsymbol{x}\in\partial\Omega\mid x_{1}=3\}$. The body force is $\boldsymbol{f}=(0,0,0)^{\top}$, and the Poisson’s ratio is fixed at $\nu=0.4$.
@@
The mechanical response of anisotropic materials (such as fiber-reinforced composites) is direction-dependent. In this scenario, the stiffness matrix $\boldsymbol{C}$ must account for the rotation of the material’s principal axes and takes the form:
$\boldsymbol{C}=\boldsymbol{T}^{\top}\hat{C}\boldsymbol{T},$ (20)
here, $\theta$ is the rotation angle of the material’s principal axis relative to the global $x$-axis, and $\boldsymbol{T}$ is the coordinate transformation matrix (Bond transformation matrix):
$\boldsymbol{T}=\begin{bmatrix}\cos^{2}\theta&\sin^{2}\theta&\cos\theta\sin\theta\\ \sin^{2}\theta&\cos^{2}\theta&-\cos\theta\sin\theta\\ -2\cos\theta\sin\theta&2\cos\theta\sin\theta&\cos^{2}\theta-\sin^{2}\theta\end{bmatrix}$, (21)
$\hat{\boldsymbol{C}}$ is the orthotropic stiffness matrix aligned with the material’s principal directions:
$\hat{\boldsymbol{C}}=\frac{1}{1-\nu_{12}\nu_{21}}\begin{bmatrix}E_{1}&\nu_{21}E_{1}&0\\ \nu_{12}E_{2}&E_{2}&0\\ 0&0&G_{12}(1-\nu_{12}\nu_{21})\end{bmatrix}$, (22)
@@
Figure 3 illustrates the computational domain. The boundary conditions are: $\Gamma_{1}=\{\boldsymbol{x}\in\partial\Omega\mid x_{1}=0\}\cup\{\boldsymbol{x}\in\partial\Omega\mid x_{2}=0\}$ is fixed, and a traction force $\boldsymbol{t}=(10^{8},\,0)^{\top}$ is applied to the upper boundary $\Gamma_{2}=\{\boldsymbol{x}\in\partial\Omega\mid x_{2}=1\}$. The body force is $\boldsymbol{f}=(0,\,0)^{\top}$.
@@
The anisotropic material parameters are sampled from the following uniform distributions (moduli in Pa):
$E_{1}\sim\mathcal{U}(50\times 10^{9},200\times 10^{9}),\quad E_{2}\sim\mathcal{U}(50\times 10^{6},200\times 10^{6}),$
$G_{12}\sim\mathcal{U}(2\times 10^{9},20\times 10^{9}),\quad\nu_{12}\sim\mathcal{U}(0.2,0.35),\quad\theta\sim\mathcal{U}(0,\pi/2).$
@@
In the 3D orthotropic case (where the material principal axes align with the global coordinate system), the stiffness matrix $\boldsymbol{C}$ is defined via the inverse of the compliance matrix:
$\boldsymbol{C}=\begin{bmatrix}\frac{1}{E_{1}}&-\frac{\nu_{21}}{E_{2}}&-\frac{\nu_{31}}{E_{3}}&0&0&0\\ -\frac{\nu_{12}}{E_{1}}&\frac{1}{E_{2}}&-\frac{\nu_{32}}{E_{3}}&0&0&0\\ -\frac{\nu_{13}}{E_{1}}&-\frac{\nu_{23}}{E_{2}}&\frac{1}{E_{3}}&0&0&0\\ 0&0&0&\frac{1}{G_{23}}&0&0\\ 0&0&0&0&\frac{1}{G_{31}}&0\\ 0&0&0&0&0&\frac{1}{G_{12}}\end{bmatrix}^{-1}.$ (23)
The parameters satisfy the reciprocity condition $\nu_{ij}E_{j}=\nu_{ji}E_{i}$. The computational domain is shown in Figure 2. The boundary conditions are: the left boundary $\Gamma_{1}$ is fixed, and a traction force $\boldsymbol{t}=(0,0,10^{8})^{\top}$ is applied to the right boundary $\Gamma_{2}$. The body force is $\boldsymbol{f}=(0,0,0)^{\top}$. The material parameters are generated as follows:
$E_{1}$ $\sim\mathcal{U}(50\times 10^{9},200\times 10^{9}),$ $E_{2}$ $\sim\mathcal{U}(50\times 10^{8},200\times 10^{8}),$ $E_{3}$ $\sim\mathcal{U}(50\times 10^{6},200\times 10^{6}),$
$G_{12}$ $\sim\mathcal{U}(2\times 10^{9},20\times 10^{9}),$ $G_{23}$ $\sim\mathcal{U}(2\times 10^{6},20\times 10^{6}),$ $G_{31}$ $\sim\mathcal{U}(2\times 10^{8},20\times 10^{8}),$
$\nu_{12}$ $\sim\mathcal{U}(0.25,0.4),$ $\nu_{13}$ $\sim\mathcal{U}(0.25,0.4),$ $\nu_{23}$ $\sim\mathcal{U}(0.25,0.4).$
@@ 4 Methods
$\mathcal{L}=\frac{1}{N_{b}}\sum_{i=1}^{N_{b}}\frac{\|\boldsymbol{f}_{i}-\boldsymbol{A}_{i}\boldsymbol{u}_{i}^{K}\|}{\|\boldsymbol{f}_{i}\|},$ (24)
where $N_{b}$ denotes the number of samples in each training batch, and $\boldsymbol{A}_{i}$ and $\boldsymbol{f}_{i}$ are the discrete system matrix and right-hand side vector corresponding to the $i$-th sample, respectively. $\boldsymbol{u}_{i}^{K}$ represents the numerical solution obtained after $K$ iterations starting from a zero initial vector using the proposed neural solves. This loss function measures the relative residual generated by substituting the model’s approximate solution into the discrete system. In practice, a balance must be struck between the number of iteration steps $K$ and the computational/memory overhead during training.
@@ 4.1 Method 1: Graph Fourier Neural Solver (G-FNS)
Figure 4: Schematic diagram of the G-FNS workflow. The upper panel depicts the hybrid iterative framework: the current iterate $\boldsymbol{u}^{k}$ is first smoothed by the operator $\mathcal{B}$ (repeated $M$ times) to obtain $\tilde{\boldsymbol{u}}^{k}$. The residual $\boldsymbol{r}^{k}$ is then processed by the neural operator $\mathcal{H}$ to predict the error correction $\boldsymbol{e}^{k}$, yielding the updated solution $\boldsymbol{u}^{k+1}$. The lower panel details the internal architecture of $\mathcal{H}$: it employs meta-networks (Meta-T and Meta-$\lambda$) that take problem parameters $\boldsymbol{\mu}$ as input to dynamically generate the transition matrices $\mathcal{C},\mathcal{C}^{*}$ and the diagonal filter $\tilde{\Lambda}$. The error $\boldsymbol{e}^{k}$ is computed through a sequence of operations involving the Fourier transform $\mathcal{F}$, the learned basis transition $\mathcal{C}$, a Hadamard product with $\tilde{\Lambda}$, and the inverse transformation.
@@
Figure 5: Schematic diagram of the Meta-network architecture (Meta-$\lambda$ or Meta-T). The input parameter field $E(\boldsymbol{x})$ is first encoded by the Left MLP Module and processed through stacked GCNConv+Linear blocks to extract spatial features. These node-level features are aggregated via Global Mean pooling into a global vector, which is subsequently mapped by the Right MLP Module and reshaped to generate the target frequency-domain operator parameters (the diagonal matrix $\tilde{\Lambda}$ or the convolution kernel of $\mathcal{C}$).
@@
  • GCNConv + Linear Module: As shown in Figure 6(a), this module feeds the input simultaneously into a GCNConv (Graph Convolutional Network) layer and a Linear layer. The outputs are summed and passed through a ReLU activation function, progressively increasing the feature dimension to $d_{l}=l\cdot d_{1},\,l=2,3,4$. The GCNConv operation (Figure 6(b)) utilizes adjacency information to aggregate neighboring node features.
@@
(a) GCNConv+Linear Module
(b) GCNConv Module
Figure 6: Schematic diagram of Meta sub-modules.
@@
We train G-FNS using Data-1. The training set size is M=12, [^0] M=12{,}000, and the computational domain is discretized using a structured mesh with $N=1089$ nodes (2178 DOFs). The smoother $\mathcal{B}$ is set to $10$ iterations of block weighted Jacobi method ($\omega=2/3$). The batch size is $64$, and the initial hidden dimension is $d_{1}=64$. The frequency domain resolution is $30\times 30$, and the number of iteration steps in the loss function is $K=5$. The Adam optimizer [43] is used with a learning rate of $1\times 10^{-4}$.
The yellow dashed line in Figure 7(a) shows the training loss trajectory. It can be observed that the loss function eventually stabilizes but converges to a relatively high value. To quantify the performance, we randomly generated 10 test samples and compared G-FNS against three variants of the SA-AMG method [44, 38]:

### v2

@@ 4 Datasets
For the 2D case, the stress tensor $\bm{\sigma}$ and strain tensor $\bm{\varepsilon}$ (using Voigt notation) are expressed as
$\bm{\sigma}=\begin{bmatrix}\sigma_{xx}\\[2.0pt] \sigma_{yy}\\[2.0pt] \sigma_{xy}\end{bmatrix},\qquad\bm{\varepsilon}=\begin{bmatrix}\varepsilon_{xx}\\[2.0pt] \varepsilon_{yy}\\[2.0pt] 2\varepsilon_{xy}\end{bmatrix}$. (32)
The corresponding isotropic stiffness matrix $\bm{C}$ takes the following form
$\bm{C}=\begin{bmatrix}\lambda+2\mu&\lambda&0\\ \lambda&\lambda+2\mu&0\\ 0&0&\mu\end{bmatrix}$,
where the Lamé constants are defined as
@@
We consider the Poisson’s ratio $\nu\in[0,0.5)$ to be constant, while the Young’s modulus $E(\bm{x})$ follows a log-normal Gaussian random field (GRF) distribution [jha2025theory] to simulate spatially heterogeneous materials
$E(\bm{x})=\alpha_{m}\cdot\exp(w(\bm{x}))+\beta_{m},$ (34)
where $\alpha_{m}=10^{8}$ and $\beta_{m}=100$. The random field $w(\bm{x})\sim\mathcal{N}(0,L_{\Delta}^{-2})$, with its covariance operator $L_{\Delta}^{-2}$ defined by the following second-order elliptic operator
@@
The computational domain is $\Omega=[0,1]^{2}$. The boundary conditions are defined as follows: the left boundary $\Gamma_{1}=\{\bm{x}\in\partial\Omega\mid x_{1}=0\}$ is fixed (Dirichlet), and a traction force $\bm{t}=(10^{6},0)^{\top}$ is applied to the right boundary $\Gamma_{2}=\{\bm{x}\in\partial\Omega\mid x_{1}=1\}$ (Neumann). The body force is set to $\bm{f}=(0,0)^{\top}$, and the Poisson’s ratio is fixed at $\nu=0.4$. Figure 2 illustrates the mesh discretization of the domain, a realization of the Young’s modulus field, and the corresponding displacement solution. By sampling the random field $E(\bm{x})$ $M$ times and discretizing the resulting PDEs, we obtain $M$ training samples.
@@
For the 3D case, the stress and strain tensors are expressed as
$\bm{\sigma}=[\sigma_{xx},\sigma_{yy},\sigma_{zz},\sigma_{yz},\sigma_{xz},\sigma_{xy}]^{\top},\quad\bm{\varepsilon}=[\varepsilon_{xx},\varepsilon_{yy},\varepsilon_{zz},2\varepsilon_{yz},2\varepsilon_{xz},2\varepsilon_{xy}]^{\top}.$ (36)
The corresponding stiffness matrix $\bm{C}$ is given by
$\bm{C}=\begin{bmatrix}\lambda+2\mu&\lambda&\lambda&0&0&0\\ \lambda&\lambda+2\mu&\lambda&0&0&0\\ \lambda&\lambda&\lambda+2\mu&0&0&0\\ 0&0&0&\mu&0&0\\ 0&0&0&0&\mu&0\\ 0&0&0&0&0&\mu\end{bmatrix}$.
The generation of Young’s modulus follows the same procedure as in Data-1. The computational domain is shown in Figure 3. The boundary conditions are: the left boundary $\Gamma_{1}=\{\bm{x}\in\partial\Omega\mid x_{1}=0\}$ is fixed, and a traction force $\bm{t}=(10^{6},0,0)^{\top}$ is applied to the right boundary $\Gamma_{2}=\{\bm{x}\in\partial\Omega\mid x_{1}=3\}$. The body force is $\bm{f}=(0,0,0)^{\top}$, and the Poisson’s ratio is fixed at $\nu=0.4$.
@@
The mechanical response of anisotropic materials (such as fiber-reinforced composites) is direction-dependent. In this scenario, the stiffness matrix $\bm{C}$ must account for the rotation of the material’s principal axes and takes the form
$\bm{C}=\bm{T}^{\top}\hat{\bm{C}}\,\bm{T},$ (37)
here, $\theta$ is the rotation angle of the material’s principal axis relative to the global $x$-axis, and $\bm{T}$ is the coordinate transformation matrix (Bond transformation matrix)
$\bm{T}=\begin{bmatrix}\cos^{2}\theta&\sin^{2}\theta&\cos\theta\sin\theta\\ \sin^{2}\theta&\cos^{2}\theta&-\cos\theta\sin\theta\\ -2\cos\theta\sin\theta&2\cos\theta\sin\theta&\cos^{2}\theta-\sin^{2}\theta\end{bmatrix}$, (38)
$\hat{\bm{C}}$ is the orthotropic stiffness matrix aligned with the material’s principal directions
$\hat{\bm{C}}=\frac{1}{1-\nu_{12}\nu_{21}}\begin{bmatrix}E_{1}&\nu_{21}E_{1}&0\\ \nu_{12}E_{2}&E_{2}&0\\ 0&0&G_{12}(1-\nu_{12}\nu_{21})\end{bmatrix}$, (39)
@@
Figure 4 illustrates the computational domain. The boundary conditions are: $\Gamma_{1}=\{\bm{x}\in\partial\Omega\mid x_{1}=0\}\cup\{\bm{x}\in\partial\Omega\mid x_{2}=0\}$ is fixed, and a traction force $\bm{t}=(10^{8},\,0)^{\top}$ is applied to the upper boundary $\Gamma_{2}=\{\bm{x}\in\partial\Omega\mid x_{2}=1\}$. The body force is $\bm{f}=(0,\,0)^{\top}$.
@@
The anisotropic material parameters are sampled from the following uniform distributions
$E_{1}\sim\mathcal{U}(50\times 10^{9},200\times 10^{9}),\quad$ $G_{12}$ $\sim\mathcal{U}(2\times 10^{9},20\times 10^{9}),$
$E_{2}\sim\mathcal{U}(50\times 10^{6},200\times 10^{6}),\quad$ $\nu_{12}$ $\sim\mathcal{U}(0.2,0.35),\quad\theta\sim\mathcal{U}(0,\pi/2).$
@@
In the 3D orthotropic case (where the material principal axes align with the global coordinate system), the stiffness matrix $\bm{C}$ is defined via the inverse of the compliance matrix
$\bm{C}=\begin{bmatrix}\frac{1}{E_{1}}&-\frac{\nu_{21}}{E_{2}}&-\frac{\nu_{31}}{E_{3}}&0&0&0\\ -\frac{\nu_{12}}{E_{1}}&\frac{1}{E_{2}}&-\frac{\nu_{32}}{E_{3}}&0&0&0\\ -\frac{\nu_{13}}{E_{1}}&-\frac{\nu_{23}}{E_{2}}&\frac{1}{E_{3}}&0&0&0\\ 0&0&0&\frac{1}{G_{23}}&0&0\\ 0&0&0&0&\frac{1}{G_{31}}&0\\ 0&0&0&0&0&\frac{1}{G_{12}}\end{bmatrix}^{-1}.$ (40)
The parameters satisfy the reciprocity condition $\nu_{ij}E_{j}=\nu_{ji}E_{i}$. The computational domain is shown in Figure 3. The boundary conditions are: the left boundary $\Gamma_{1}$ is fixed, and a traction force $\bm{t}=(0,0,10^{8})^{\top}$ is applied to the right boundary $\Gamma_{2}$. The body force is $\bm{f}=(0,0,0)^{\top}$. The material parameters are generated as follows
$E_{1}$ $\sim\mathcal{U}(50\times 10^{9},200\times 10^{9}),$ $G_{12}$ $\sim\mathcal{U}(2\times 10^{9},20\times 10^{9}),$ $\nu_{12}$ $\sim\mathcal{U}(0.25,0.4),$
$E_{2}$ $\sim\mathcal{U}(50\times 10^{8},200\times 10^{8}),$ $G_{23}$ $\sim\mathcal{U}(2\times 10^{6},20\times 10^{6}),$ $\nu_{13}$ $\sim\mathcal{U}(0.25,0.4),$
$E_{3}$ $\sim\mathcal{U}(50\times 10^{6},200\times 10^{6}),$ $G_{31}$ $\sim\mathcal{U}(2\times 10^{8},20\times 10^{8}),$ $\nu_{23}$ $\sim\mathcal{U}(0.25,0.4).$
@@ 5 Methods
$\mathcal{L}=\frac{1}{N_{b}}\sum_{i=1}^{N_{b}}\frac{\|\bm{f}_{i}-\bm{A}_{i}\bm{u}_{i}^{(K)}\|}{\|\bm{f}_{i}\|},$ (41)
where $N_{b}$ denotes the number of samples in each training batch, and $\bm{A}_{i}$ and $\bm{f}_{i}$ are the discrete system matrix and right-hand side vector corresponding to the $i$-th sample, respectively. $\bm{u}_{i}^{(K)}$ represents the numerical solution obtained after $K$ iterations starting from a zero initial vector using the proposed neural solves. This loss function measures the relative residual generated by substituting the model’s approximate solution into the discrete system. In practice, a balance must be struck between the number of iteration steps $K$ and the computational/memory overhead during training.
@@ 5.1 Method 1: Graph Fourier Neural Solver (G-FNS)
Figure 5: Schematic diagram of the G-FNS workflow. The upper panel depicts the hybrid iterative framework: the current iterate $\bm{u}^{(k)}$ is first smoothed by the operator $\mathcal{B}$ (repeated $M$ times) to obtain $\bm{u}^{(k+\frac{1}{2})}$. The residual $\bm{r}^{(k+\frac{1}{2})}$ is then processed by the neural operator $\mathcal{H}$ to predict the error correction $\bm{e}^{(k+\frac{1}{2})}$, yielding the updated solution $\bm{u}^{(k+1)}$. The lower panel details the internal architecture of $\mathcal{H}$: it employs meta-networks (Meta-T and Meta-$\lambda$) that take problem parameters $\bm{\mu}$ as input to dynamically generate the transition matrices $\mathcal{C},\mathcal{C}^{*}$ and the diagonal filter $\tilde{\Lambda}$. The error $\bm{e}^{(k+\frac{1}{2})}$ is computed through a sequence of operations involving the Fourier transform $\mathcal{F}$, the learned basis transition $\mathcal{C}$, a Hadamard product with $\tilde{\Lambda}$, and the inverse transformation.
@@
Figure 6: Schematic diagram of the Meta-network architecture (Meta-$\lambda$ or Meta-T). The input parameter field $E(\bm{x})$ is first encoded by the Left MLP Module and processed through stacked GCNConv+Linear blocks to extract spatial features. These node-level features are aggregated via Global Mean pooling into a global vector, which is subsequently mapped by the Right MLP Module and reshaped to generate the target frequency-domain operator parameters (the diagonal matrix $\tilde{\Lambda}$ or the convolution kernel of $\mathcal{C}$).
@@
  • GCNConv + Linear Module: This module feeds the input simultaneously into a GCNConv[kipf2017semi] (Graph Convolutional Network) layer and a Linear layer. The outputs are summed and passed through a ReLU activation function, progressively increasing the feature dimension to $d_{l}=l\cdot d_{1},\,l=2,3,4$.
@@
We train G-FNS using Data-1. The training set size is M=12, [^0] M=12{,}000, and the computational domain is discretized using a structured mesh with $N=1089$ nodes (2178 DOFs). The smoother $\mathcal{B}$ is set to $10$ iterations of weighted block Jacobi Method ($\omega=2/3$). The batch size is $64$, and the initial hidden dimension is $d_{1}=64$. The frequency domain resolution is $30\times 30$, and the number of iteration steps in the loss function is $K=5$. The Adam optimizer [adam2014method] is used with a learning rate of $1\times 10^{-4}$.
The yellow dashed line in Figure 7(a) shows the training loss trajectory. It can be observed that the loss function eventually stabilizes but converges to a relatively high value. To quantify the performance, we randomly generated 10 test samples and compared G-FNS against three variants of the SA-AMG method [vanvek1996algebraic, EnergyMinimization2011]: